# Nonequilibrium entanglement between levitated masses under optimal control

### v1

@@ Abstract
In this work, we combine methods of nonequilibrium entanglement generation with the time-continuous quantum control [27] of two harmonically trapped masses interacting in free space. We leverage continuous measurement and feedback with a periodic modulation of the interaction to drive the system into a non-equilibrium state that exhibits strong unconditional entanglement.We show how an optimal quantum measurement can be performed to generate entanglement in the conditional state [28]. Following this strategy, we connect the observable entanglement to the properties of the measurement apparatus [27]. Then we use feedback control to stabilize the effective temperature of the motion, thus optimizing entanglement likelihood through a specifically chosen cost function. Ultimately, our protocol achieves unconditional (out-of-loop) entanglement between the motion of directly interacting masses at the fundamental limit of their joint conditional state.
@@
Figure 1: Conditional period-averaged logarithmic negativity as the function of driving amplitude $g_{1}$ and modulation frequency $\Omega$. Panels on the right illustrate the standard deviation of the bivariate Wigner distribution, given by the conditional covariance matrix $\Sigma^{c}$, depending on $X_{1}$ and $P_{1}$ for three characteristic values of $\{g_{1}/\Omega_{0},\Omega/\Omega_{0}\}=\{0.17,2.7\}$,$\{0.08,2.5\}$,$\{0.01,2.0\}$, labeled (i–iii) on the color plot. Here we use the following parameters of the system: $g_{0}/\Omega_{0}=0.2$, $\Gamma_{\rm ba}/\Omega_{0}=5\%$, $\Gamma_{\rm th}/\Gamma_{\rm ba}=5\%$ and $q=0.1/\Omega_{0}$, fully consistent with the parameters in Ref. [28].
@@
We consider two identical particles with mass $m$, each confined in one spatial dimension by a harmonic potential with eigenfrequency $\Omega_{0}$ as in Ref.  [28]. The particles are modeled to interact via a central potential of the form $k/r^{n}$, e.g. via gravitation, electrostatic or Casimir interaction, where we introduce $r=\sqrt{x_{\rm zpf}^{2}\left(X_{1}-X_{2}\right)^{2}+R^{2}}$ with particle positions $X_{k}$. Here, $R$ denotes the interparticle distance along a spatial direction orthogonal to the motion $X_{1,2}$ with $x_{\rm zpf}=\sqrt{\hbar/m\Omega_{0}}$. In the limit of small displacements and considering terms up to quadratic order in position yields the interaction Hamiltonian $H_{\rm int}=\hbar g(X_{1}-X_{2})^{2}$ with a coupling rate $g=k/(2nR^{2+n}m\Omega_{0})$. For the following discussion, we assume that the coupling rate can be temporally modulated with the frequency $\Omega$ e.g. by varying the distance $R$, such that $g(t)=g_{0}+2g_{1}\cos\Omega t$. Additionally, the particles are controlled individually by a feedback force $u_{k}$. The total Hamiltonian is then given by
@@
In the rotating reference frame, the parametric modulation of the coupling $g(t)$ leads to the term $\propto(a_{1}^{\dagger}a_{2}^{\dagger}-a_{1}a_{2})$ [29, 30], where $a_{1,2}$ are the annihilation operators for the oscillator motion. This term describes two-mode squeezing and entanglement generation.
Relevant decoherence channels include thermal decoherence from residual gas and photon recoil with rates $\Gamma_{\rm th}$ and $\Gamma_{\rm ba}$ respectively. These settings are typical for various continuously measured interferometric arrangements, such as optical trapping of sub-micron particles in ultra-high vacuum, which offers extreme isolation, scalability, and potential for landscape engineering [3].
@@
$+\Sigma^{c}(t)\mathbf{C}^{T}\mathbf{W}^{-1}d\mathbf{w}(t)$ (3a)
@@
with control matrix $\mathbf{B}$, input force vector $\mathbf{u}=(u_{1},u_{2})^{T}$, measurement matrix $\mathbf{C}$, system and measurement noise correlation matrices $\mathbf{V}$ and $\mathbf{W}$ and stochastic vector $d\mathbf{w}(t)=\left(dW_{1}(t),dW_{2}(t)\right)^{T}$.
Making the strong interferometric measurement of the position quadrature unavoidably perturbs the system, causing it to explore a larger phase space. Subject to excess noise, the averaged conditional state turns into the unconditional state of the system $\rho=\mathbb{E}[\rho_{c}]$ [27]. The covariance matrix of the unconditional state is given by the sum of the conditional state and the excess noise $\Sigma^{u}=\Sigma^{c}+{\Xi^{ex}}$.
@@
The optimal feedback $\mathbf{u}(t)=-\mathbf{K}(t)\mathbf{X}_{c}(t)$ is determined by the feedback gain $\mathbf{K}(t)=\mathbf{B}^{T}\Omega(t)/q$, solving the backwards Riccati equation
$d\Omega(t)=-\mathbf{A}^{T}(t)\Omega(t)-\Omega(t)\mathbf{A}(t)dt-\mathbf{P}dt+\Omega(t)\frac{\mathbf{B}\mathbf{B}^{T}}{q}\Omega(t)dt.$ (4)
The excess noise matrix ${\Xi^{ex}}$ depends on both, the measurement setup and the feedback from the controller, it’s time evolution is governed by  [27]
$\begin{gathered}d{\Xi^{ex}}(t)=\left(\mathbf{A}-\mathbf{B}\mathbf{K}(t)\right){\Xi^{ex}}(t)dt\\ +{\Xi^{ex}}(t)\left(\mathbf{A}-\mathbf{B}\mathbf{K}(t)\right)^{T}dt\\ +\left(\Sigma^{c}(t)\mathbf{C}^{T}+\mathbf{M}\right)\mathbf{W}^{-1}\left(\Sigma^{c}(t)\mathbf{C}^{T}+\mathbf{M}\right)^{T}dt.\end{gathered}$ (5)
@@
Figure 2: The boundary for the conditional and unconditional time-averaged entanglement is shown by the blue and dark red surfaces, respectively. The case of attractive interaction $g_{0}>0$ (a) and repulsive interaction $g_{0}<0$ (b). All calculations are done for the EPR phase $\varphi=\pi$. The contour curves (solid) show the boundaries for unconditional entanglement calculated at $\eta=0.25,~{}0.5,~{}0.75$ and $1$, with their projections illustrated by the dotted curves.
@@
Figure 3: Time-dependent logarithmic negativity $\mathcal{E}_{\mathcal{N}}(t)$ for (a) attractive case with $g_{0}/\Omega_{0}=0.2$ and (b) repulsive case with $g_{0}/\Omega_{0}=-0.2$. Dotted black curves correspond to the conditional negativity, solid black curves correspond to the unconditional one. Horizontal solid lines show the values of $\mathcal{E}_{\mathcal{N}}(t)$ without the modulation, $g_{1}=0$. Calculated for $\Omega/\Omega_{0}=2.74$ (a) and $\Omega/\Omega_{0}=0.7$ (b).
@@
Figure 1 shows the conditional entanglement measure, logarithmic negativity $\langle\mathcal{E}_{N}^{c}\rangle$, calculated as a function of the amplitude $g_{1}$ and frequency $\Omega$ of the modulation. Red color regions correspond to the entangled conditional state, where $\langle\mathcal{E}_{N}^{c}\rangle>0$. This happens at the parametric resonances, with the main resonant frequency being $\Omega=2\Omega^{*}$, $2\Omega^{*}=2\sqrt{\Omega_{0}^{2}+4\Omega_{0}g_{0}}\approx 2.74\Omega_{0}$ for the chosen parameters. When the squeezing in the joint phase space is strong enough quantum correlations survive thermal fluctuations destroying it otherwise. Increasing the modulation depth $g_{1}$ expands the entangled region. Furthermore, for sufficiently large $g_{1}$, an additional parametric resonance emerges at the half of the drive frequency $\Omega=\Omega^{*}$. This result is well-known from the stability analysis of quantum systems under parametric drives [45]. The two-mode squeezing can also be achieved through parametric driving of the particle frequencies using the correct phase relation ${\varphi}$ between the frequency drives $\Omega_{m}(t)=\Omega_{0}+\Omega_{1}\cos(\Omega_{p}t+{\varphi})$, Here, a modulation of the interaction strength naturally ensures the perfect phases for optimal squeezing, which is evidenced in Fig. 1(i-iii). The panels (i–iii) illustrate the standard deviation of the corresponding Wigner function, characterized by the covariance matrix $\Sigma^{c}$ of the conditional state, calculated for several characteristic values of $\{g_{1},\Omega_{1}\}$.
@@
Having established conditional entanglement generation via parametric modulation, we now explore a broader parameter space. The key parameters are the stationary interaction strength $g_{0}/\Omega_{0}$, the amplitude $g_{1}/g_{0}$ and frequency $\Omega/\Omega_{0}$ of the time-dependent interaction, and the detection efficiency $\eta$ of the time-continuous position measurement. To present the calculated $\langle\mathcal{E}_{N}(g_{0},g_{1},\eta,\Omega)\rangle$ within a single plot, for every combination of $\{g_{0},g_{1},\eta\}$, we select the maximum logarithmic negativity at the frequency $\Omega$, corresponding to the strongest parametric resonance in the system. Figure 2 shows the separability boundaries $\langle\mathcal{E}_{N}^{c}\rangle=0$ obtained in the case of attractive (a) and repulsive (b) interaction. Here, we represent the performance of both conditional (blue surface) and unconditional states (red surface). We also plot contours on the surface of the conditional state for $\eta$ equal to 0.25, 0.5, 0.75, and 1, along with projections on the $(g_{0},g_{1})$ plane. In the attractive regime with relatively low modulation depth $g_{1}=0.2g_{0}$ and realistic detection efficiency of $\eta=0.5$, our approach predicts entanglement generation at $g_{0}=0.1\Omega_{0}$. This is more than an order of magnitude lower in comparison to existing stationary protocols [18, 28]. Stationary entanglement in the repulsive case requires $|g_{0}|>0.2\Omega_{0}$ [28], but the system turns unstable when $|g_{0}|\geq 0.25\Omega_{0}$, which makes it extremely hard to implement. The parametric drive of the interaction modifies fundamental limits of entanglement generation, making it possible under more relaxed experimental conditions, Fig. 2b.
@@
Figure 3 demonstrates the time dependence of the logarithmic negativity of unconditional states (black solid curves) for the attractive (a) and repulsive (b) regimes. For comparison, we also include the results without the modulation (red dotted lines) and the negativity of conditional states with modulation (black dotted curves). We observe a significant increase in the logarithmic negativity of the unconditional state, surpassing the separability boundary $\mathcal{E}_{N}^{u}=0$, and approaching the fundamental limit of nonequilibrium entanglement. This is a remarkable result that allows for out-of-loop entanglement generation between large masses under realistic experimental conditions. Importantly, the peaking period of the unconditional entanglement is persistent with the period of parametric resonance $1/\Omega$. Similar phenomena have been observed in amplitude-modulated optomechanical systems, where the entanglement between the mirror motion and the cavity field oscillates with the period of the parametric drive [25].

### v2

@@ Abstract
Figure 1: Schematic illustration of the non-equilibrium entanglement generation between two levitating masses $m$, interacting via $\sim{1/r^{n}}$ potential in free space, subject to parametric modulation of the separation R(t). Both masses undergo continuous position measurement (Det) and feedback control (LQR) steering their motion to the entangled state.
@@
In this work, we combine methods of nonequilibrium entanglement generation with the time-continuous quantum control [27] of two harmonically trapped masses interacting in free space. Figure 1 presents a schematic illustration of the system under consideration. We leverage continuous measurement and feedback with a periodic modulation of the interaction to drive the system into a non-equilibrium state that exhibits strong unconditional entanglement.We show how an optimal quantum measurement can be performed to generate entanglement in the conditional state [28]. Following this strategy, we connect the observable entanglement to the properties of the measurement apparatus [27]. Then we use feedback control to stabilize the effective temperature of the motion, thus optimizing entanglement likelihood through a specifically chosen cost function. Ultimately, our protocol achieves unconditional (out-of-loop) entanglement between the motion of directly interacting masses at the limit of their joint conditional state.
@@
Figure 2: Conditional period-averaged logarithmic negativity as the function of driving amplitude $g_{1}$ and modulation frequency $\Omega$. Panels on the right illustrate the standard deviation of the bivariate Wigner distribution, given by the conditional covariance matrix $\Sigma^{c}$, depending on coordinates and momenta of the symmetric (common) and antisymmetric (differential) modes $X_{1}\pm X_{2}$ and $P_{1}\pm P_{2}$ for a characteristic value of $\{g_{1}/g_{0},\Omega/\Omega_{0}\}=\{0.3,2.7\}$, labeled by a star on the color plot. Here we use the following parameters of the system: $g_{0}/\Omega_{0}=0.2$, $\Gamma_{\rm ba}/\Omega_{0}=5\%$, $\Gamma_{\rm th}/\Gamma_{\rm ba}=5\%$ and $q=0.1/\Omega_{0}$, fully consistent with the parameters in Ref. [28].
@@
We consider two identical particles with mass $m$, each confined in one spatial dimension by a harmonic potential with eigenfrequency $\Omega_{0}$ as shown in Fig. 1. The particles are modeled to interact via a central potential of the form $k/r^{n}$, e.g. via gravitation, electrostatic or Casimir interaction, where we introduce $r=\sqrt{x_{\rm zpf}^{2}\left(X_{1}-X_{2}\right)^{2}+R^{2}}$ with particle positions $X_{k}$. Here, $R$ denotes the interparticle distance along a spatial direction orthogonal to the motion $X_{1,2}$ with $x_{\rm zpf}=\sqrt{\hbar/m\Omega_{0}}$. In the limit of small displacements and considering terms up to quadratic order in position yields the interaction Hamiltonian $H_{\rm int}=\hbar g(X_{1}-X_{2})^{2}$ with a coupling rate $g=k/(2nR^{2+n}m\Omega_{0})$. For the following discussion, we assume that the coupling rate can be temporally modulated with the frequency $\Omega$ e.g. by varying the distance $R$, such that $g(t)=g_{0}+2g_{1}\cos\Omega t$. Additionally, the particles are controlled individually by a feedback force $u_{k}$. The total Hamiltonian is then given by
@@
The interaction term acts selectively on the differential mode leading to quadrature squeezing $\sim\hbar g(t)\left(\hat{a}_{-}^{2}+\hat{a}_{-}^{\dagger 2}\right)$, where $\hat{a}_{-}=(\hat{a}_{1}-\hat{a}_{2})/\sqrt{2}$ is the differential-mode annihilation operator, defined via the bare ladder operators $\hat{a}_{1}$ and $\hat{a}_{2}$ of the two oscillators. When the time-dependent interaction strength hits the resonant frequency $\Omega=2\Omega_{-}$, the squeezing is further enhanced $\sim g_{1}$, leading to the buildup of strong non-classical correlations in the out-of-phase motion of the particles and resulting in entanglement generation (see Supplementary Materials for detailed analysis).
Relevant decoherence channels include thermal decoherence from residual gas and photon recoil with rates $\Gamma_{\rm th}=\gamma n_{\rm th}$ [7] and $\Gamma_{\rm ba}$ respectively. Here, the mean thermal occupation number $n_{\rm th}$ quantifies the thermal energy of the environments each particle is in contact with, and is given in the high-temperature limit by $n_{\rm th}\approx k_{\rm B}T/{\hbar\Omega_{0}}$. These settings are typical for various continuously measured interferometric arrangements, such as optical trapping of sub-micron particles in ultra-high vacuum, which offers extreme isolation, scalability, and potential for landscape engineering [3].
@@
$+\Sigma^{c}(t)\mathbf{C}^{T}\mathbf{W}^{-1}d\mathbf{w}(t)\>,$ (3a)
@@
with control matrix $\mathbf{B}$, input force vector $\mathbf{u}=(u_{1},u_{2})^{T}$, measurement matrix $\mathbf{C}$, system and measurement noise correlation matrices $\mathbf{V}$ and $\mathbf{W}$ and stochastic vector $d\mathbf{w}(t)=\left(dW_{1}(t),dW_{2}(t)\right)^{T}$. All presented matrices are displayed in their full form in the end matter of this work.
Making the interferometric measurement of the position quadrature unavoidably perturbs the system, causing it to explore a larger phase space. Subject to excess noise, the averaged conditional state turns into the unconditional state of the system $\rho=\mathbb{E}[\rho_{c}]$ [27]. The covariance matrix of the unconditional state is given by the sum of the conditional state and the excess noise $\Sigma^{u}=\Sigma^{c}+{\Xi^{ex}}$. Since ${\Xi^{ex}},\Sigma^{c}$ are positive definite, the conditional state occupies a smaller region in phase space compared to the unconditional state. As the conditional state is fully determined by the measurement setup and independent of feedback, it imposes a fundamental limit on the entanglement of the unconditional state for a given finite interaction rate[28].
@@
From now on, we will use the Einstein-Podolsky-Rosen (EPR) cost function, which minimizes EPR-type variances in the joint phase space of the two particles, as detailed in Ref. [28]. The optimal feedback $\mathbf{u}(t)=-\mathbf{K}(t)\mathbf{X}_{c}(t)$ is determined by the feedback gain $\mathbf{K}(t)=\mathbf{B}^{T}\Omega(t)/q$, solving the backwards Riccati equation
$d\Omega(t)=-[\mathbf{A}^{T}(t)\Omega(t)+\Omega(t)\mathbf{A}(t)+\mathbf{P}]dt+\Omega(t)\frac{\mathbf{B}\mathbf{B}^{T}}{q}\Omega(t)dt.$ (4)
The excess noise matrix ${\Xi^{ex}}$ depends on both, the measurement setup and the feedback from the controller, its time evolution is governed by  [27]
$d{\Xi^{ex}}(t)=$ $\left(\mathbf{A}-\mathbf{B}\mathbf{K}(t)\right){\Xi^{ex}}(t)dt$ (5)
$+$ ${\Xi^{ex}}(t)\left(\mathbf{A}-\mathbf{B}\mathbf{K}(t)\right)^{T}dt$
$+$ $\left(\Sigma^{c}(t)\mathbf{C}^{T}+\mathbf{M}\right)\mathbf{W}^{-1}\left(\Sigma^{c}(t)\mathbf{C}^{T}+\mathbf{M}\right)^{T}dt\>.$
@@
Figure 3: The boundary for the conditional and unconditional time-averaged entanglement is shown by the blue and dark red surfaces, respectively. The case of attractive interaction $g_{0}>0$ (a) and repulsive interaction $g_{0}<0$ (b). Calculation has been performed for the EPR phase $\theta=\pi$ (a) and $\theta=0$ (b) . The contour curves (solid) show the boundaries for unconditional entanglement calculated at $\eta=0.25,~{}0.5,~{}0.75$ and $1$, with their projections illustrated by the dotted curves.
@@
Figure 4: Time-dependent logarithmic negativity $\mathcal{E}_{\mathcal{N}}(t)$ at detection efficiency $\eta=0.5$ for (a) attractive case with $g_{0}/\Omega_{0}=0.2$ and (b) repulsive case with $g_{0}/\Omega_{0}=-0.2$. Dotted black curves correspond to the conditional negativity, solid black curves correspond to the unconditional one. Horizontal solid lines show the values of $\mathcal{E}_{\mathcal{N}}(t)$ without the modulation, $g_{1}=0$. Calculated for $|g_{1}/g_{0}|=0.15$, $\Omega/\Omega_{0}=2.74$, EPR phase $\theta=\pi$ (a) and $\Omega/\Omega_{0}=0.7$, $\theta=0$ (b).
@@
Figure 2 shows the conditional entanglement measure, logarithmic negativity $\langle\mathcal{E}_{N}^{c}\rangle$, calculated as a function of the amplitude $g_{1}$ and frequency $\Omega$ of the modulation. Red color regions correspond to the entangled conditional state, where $\langle\mathcal{E}_{N}^{c}\rangle>0$. This happens at the parametric resonances, with the main resonant frequency being $\Omega=2\Omega^{*}$, $2\Omega^{*}=2\sqrt{\Omega_{0}^{2}+4\Omega_{0}g_{0}}\approx 2.74\Omega_{0}$ for the chosen parameters. When the squeezing in the joint phase space is strong enough quantum correlations survive thermal fluctuations destroying it otherwise. Increasing the modulation depth $g_{1}$ expands the entangled region. Furthermore, for sufficiently large $g_{1}$, an additional parametric resonance emerges at the half of the drive frequency $\Omega=\Omega^{*}$. This result is well-known from the stability analysis of quantum systems under parametric drives [46]. The two-mode squeezing can also be achieved through parametric driving of the particle frequencies using the correct phase relation ${\varphi}$ between the frequency drives $\Omega_{m}(t)=\Omega_{0}+\Omega_{1}\cos(\Omega_{p}t+{\varphi})$. In the end matter of this work, we also explore the potential of this parametric driving, discussing the results and the prospects of combining this approach with our current strategy of temporal coupling modulation to enhance the overall entanglement generation. However, a modulation of the interaction strength naturally ensures the perfect phases for optimal squeezing, which is evidenced in Fig. 2(i-iii). The panels (i–iii) illustrate the standard deviation of the corresponding Wigner function, characterized by the covariance matrix $\Sigma^{c}$ of the conditional state, calculated for the symmetric and anti-symmetric state with position and momentum quadrature $X_{1}\pm X_{2}$ and $P_{1}\pm P_{2}$ respectively as well as the Wigner function of $X_{1}$ versus $X_{2}$.
@@
Having established conditional entanglement generation via parametric modulation, we now explore a broader parameter space. The key parameters are the stationary interaction strength $g_{0}/\Omega_{0}$, the amplitude $g_{1}/g_{0}$ and frequency $\Omega/\Omega_{0}$ of the time-dependent interaction, and the detection efficiency $\eta$ of the time-continuous position measurement. To present the calculated $\langle\mathcal{E}_{N}(g_{0},g_{1},\eta,\Omega)\rangle$ within a single plot, for every combination of $\{g_{0},g_{1},\eta\}$, we select the maximum logarithmic negativity at the frequency $\Omega$, corresponding to the strongest parametric resonance in the system. Figure 3 shows the separability boundaries $\langle\mathcal{E}_{N}^{c}\rangle=0$ obtained in the case of attractive (a) and repulsive (b) interaction. Here, we represent the performance of both conditional (blue surface) and unconditional states (red surface). We also plot contours on the surface of the conditional state for $\eta$ equal to 0.25, 0.5, 0.75, and 1, along with projections on the $(g_{0},g_{1})$ plane. In the attractive regime with relatively low modulation depth $g_{1}=0.2g_{0}$ and realistic detection efficiency of $\eta=0.5$, our approach predicts entanglement generation at $g_{0}=0.1\Omega_{0}$. This is more than an order of magnitude lower in comparison to existing stationary protocols [18, 28]. Stationary entanglement in the repulsive case requires $|g_{0}|>0.2\Omega_{0}$ [28], but the system turns unstable when $|g_{0}|\geq 0.25\Omega_{0}$, which makes it extremely hard to implement. The parametric drive of the interaction modifies limits of entanglement generation fundamentally set by the conditional state, making it possible under more relaxed experimental conditions, Fig. 3b.
@@
Figure 4 demonstrates the time dependence of the logarithmic negativity of unconditional states (black solid curves) for the attractive (a) and repulsive (b) regimes. For comparison, we also include the results without the modulation (red dotted lines) and the negativity of conditional states with modulation (black dotted curves). We observe a significant increase in the logarithmic negativity of the unconditional state, surpassing the separability boundary $\mathcal{E}_{N}^{u}=0$, and approaching the fundamental limit of nonequilibrium entanglement expected for the conditional state. This is a remarkable result that allows for out-of-loop entanglement generation between large masses under realistic experimental conditions. Importantly, the peaking period of the unconditional entanglement is persistent with the period of parametric resonance $1/\Omega$. Similar phenomena have been observed in amplitude-modulated optomechanical systems, where the entanglement between the mirror motion and the cavity field oscillates with the period of the parametric drive [25].